# Identification of Dynamic Nonlinear Panel Models

### v2

@@ A.6 Proof of Lemma 1
When all regressors are exogenous in the sense $\epsilon_{is}\mid(W_{ist},\alpha_{i})\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid(W_{ist},\alpha_{i})$, the identifying restriction in Proposition 4 is given as
$1-\sum_{j=1}^{J}\Pr(Y_{is}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j}-w_{s}^{\prime}\theta_{0}\geq a\}\\ \geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}.$ (19)
For any $a\in\mathcal{R}$, the above identifying restriction is only informative when there exists $2\leq j_{1}\leq J$ and $11\leq j_{2}\leq J-$ 1 such that
$b_{j_{1}}-w_{s}^{\prime}\theta_{0}\geq a,\quad b_{j_{2}+1}-w_{t}^{\prime}\theta_{0}\leq a.$ (20)
It is because if the above condition does not hold, then either the upper bound in condition (19) becomes one or the lower bound becomes zero. Then condition (19) holds for any $\theta$ and there is no identifying power.
Condition (20) can be satisfied for some $a$ if and only if there exists $2\leq j_{1}\leq J$ and $11\leq j_{2}\leq J-$ 1 such that $b_{j_{1}}-w_{s}^{\prime}\theta_{0}\geq b_{j_{2}+1}-w_{t}^{\prime}\theta_{0}$. In this case, the identifying condition (19) becomes
if $\text{if }b_{j_{1}}-w_{s}^{\prime}\theta_{0}\geq b_{j_{2}+1}-w_{t}^{\prime}\theta_{0}\Longrightarrow\sum_{j=1}^{j_{1}-1}\Pr(Y_{is}=y_{j}\mid w_{st})=1-\sum_{j=j_{1}}^{J}\Pr(Y_{is}=y_{j}\mid w_{st})\\ \geq\sum_{j=1}^{j_{2}}\Pr(Y_{it}=y_{j}\mid w_{st}).$
Replacing $j_{1}$ with $\tilde{j}_{1}=j_{1}-1$ yielding the results in Proposition 4.
A.7 Proof of Proposition 5 and Proposition 6
Proof.
We first prove that $\beta_{0}$ is point identified up to scale by showing that for any $\Delta z\in\Delta\mathcal{Z}_{order}$, the sign of $\Delta z^{\prime}\beta_{0}$ is identified. The identifying condition for $\theta_{0}$ in Proposition 4 is given as
$1-\sum_{j=1}^{J}\Pr(Y_{i1}=y_{j},b_{j}-z_{1}^{\prime}\beta_{0}-X_{i1}^{\prime}\gamma_{0}\geq a\mid z_{12})\geq\sum_{j=1}^{J}\Pr(Y_{i2}=y_{j},b_{j+1}-z_{2}^{\prime}\beta_{0}-X_{i2}^{\prime}\gamma_{0}\leq a\mid z_{12}),$
and the same condition also holds when changing the order of period 1 and 2.
Let $a=b_{k}-z_{1}^{\prime}\beta_{0}-x^{\prime}\gamma_{0}$ for $2\leq k\leq J$, then we know that $b_{j}-z_{1}^{\prime}\beta_{0}-x^{\prime}\gamma_{0}\geq a$ if $j\geq k$. Therefore, the upper bound becomes
$1-\sum_{j=1}^{J}\Pr(Y_{i1}=y_{j},b_{j}-z_{1}^{\prime}\beta_{0}-X_{i1}^{\prime}\gamma_{0}\geq a\mid z_{st})\geq 1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x).$
Moreover, if $-z_{2}^{\prime}\beta_{0}\leq-z_{1}^{\prime}\beta_{0}$, then $b_{j+1}-z_{2}^{\prime}\beta_{0}-x^{\prime}\gamma_{0}\leq a$ for $j+1\leq k$. The lower bound can be bounded below by
$\sum_{j=1}^{J}\Pr(Y_{i2}=y_{j},b_{j+1}-z_{2}^{\prime}\beta_{0}-X_{i2}^{\prime}\gamma_{0}\leq a\mid z_{12})\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12}).$
Combining the results yields the following conditions:
$\Delta z^{\prime}\beta_{0}\geq 0\Longrightarrow 1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x)\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12}).$
The contraposition of the above condition generates
$01-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x)\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12})\Longrightarrow\Delta z^{\prime}\beta_{0}<$ 0 .
Therefore, when $z\in\mathcal{Z}_{1,order}$, the sign of $\Delta z^{\prime}\beta_{0}<0$ is identified. Similarly, the sign of $\Delta z^{\prime}\beta_{0}>0$ is also identified when $z\in\mathcal{Z}_{2,order}$. Given that the sign of $\Delta z^{\prime}\beta_{0}$ is identified, the parameter $\beta_{0}$ is point identified up to scale under the large support condition $\Delta z$. The analysis is the same as proof A.3 for Proposition 2, so it is omitted here.
Now we look at the result for $\gamma_{0}$ in Proposition 6. Let $a=b_{k}-z_{1}^{\prime}\beta_{0}-x_{1}^{\prime}\gamma_{0}$, then the upper bound becomes
$1-\sum_{j=1}^{J}\Pr(Y_{i1}=y_{j},b_{j}-z_{1}^{\prime}\beta_{0}-X_{i1}^{\prime}\gamma_{0}\geq a\mid z_{st})\geq 1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1}).$
If $b_{k}-z_{2}^{\prime}\beta_{0}-x_{2}^{\prime}\gamma_{0}\leq a$, the lower bound becomes
$\sum_{j=1}^{J}\Pr(Y_{i2}=y_{j},b_{j+1}-z_{2}^{\prime}\beta_{0}-X_{i2}^{\prime}\gamma_{0}\leq a\mid z_{12})\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x_{2}\mid z_{12}).$
Combining the results leads to
if $\text{if }-z_{2}^{\prime}\beta_{0}-x_{2}^{\prime}\gamma_{0}\leq-z_{1}^{\prime}\beta_{0}-x_{1}^{\prime}\gamma_{0}\Longrightarrow\\ 1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1})\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12}).$
Similarly, the contraposition of the above condition gives
$1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1})<\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x_{2}\mid z_{12})\Longrightarrow\Delta z^{\prime}\beta_{0}<(x_{1}-x_{2})^{\prime}\gamma_{0}.$
Therefore, when $\Delta z\in\mathcal{Z}^{j}_{3,order}\cap\mathcal{Z}_{2,order}$, it implies that
$0<\Delta z^{\prime}\beta_{0}<(x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}.$
Given that $x_{1}^{j}-x_{2}^{j}\neq 0$, it implies that the sign of $\gamma^{j}_{0}$ is identified. Similarly when $\Delta z\in\mathcal{Z}^{j}_{4,order}\cap\mathcal{Z}_{1,order}$, we have $0<\Delta z^{\prime}\beta_{0}<(x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}$. The sign of $\gamma^{j}_{0}$ is also identified. For any $j\leq d_{x}$, Proposition 6 requires that either $\mathcal{Z}^{j}_{3,order}\cap\mathcal{Z}_{2,order}\neq\emptyset$ or $\mathcal{Z}^{j}_{4,order}\cap\mathcal{Z}_{1,order}\neq\emptyset$, so the sign of $\gamma^{j}_{0}$ is identified for any $j$.
A.8 Proof of Proposition 7
Proof.
For any set $K\subset\mathcal{J}$, the conditional probability of selecting a choice $j\in K$ given $W_{ist}=w_{st}$ is:
$\Pr(Y_{it}\in K\mid w_{st})=\Pr\left(\exists j\in K\text{ s.t. }w_{ijt}^{\prime}\theta_{0}+v_{ijt}\geq w_{ikt}^{\prime}\theta_{0}+v_{ikt}\ \forall k\in K^{c}\mid w_{st}\right).$
The above observed probability provides information regarding the distribution of $v_{ikt}-v_{ijt}$. Therefore, we can use observed data to bound the following distribution: for $a_{jk}\in\mathcal{R}$,
$Q_{t}(a_{jk}\mid w_{st}):=\Pr\left(\exists j\in K\text{ s.t. }v_{ikt}-v_{ijt}\leq a_{jk}\ \forall k\in K^{c}\mid w_{st}\right).$
When $a_{jk}$ satisfies $a_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ for any $(j,k)$, then the above probability can be bounded below as
$Q_{t}(a_{jk}\mid w_{st})$ $\geq\Pr\left(\exists j\in K\text{ s.t. }v_{ikt}-v_{ijt}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall k\in K^{c}\mid w_{st}\right)$
$=\Pr(Y_{it}\in K\mid w_{st}).$
Therefore, the lower bound for $Q_{t}(a_{jk}\mid w_{st})$ is established as
$Q_{t}(a_{jk}\mid w_{st})\geq\Pr(Y_{it}\in K\mid w_{st})\mathbf{\mathbbm{1}}\{a_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,k\in K^{c}\}.$
The above inequality holds since either $a_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ or the lower bound is zero.
Since Assumption 1 is only conditioning on the exogenous covariate $z_{st}$, we can bound the conditional distribution $Q_{t}(a_{jk}\mid z_{st})$ by taking expectation of $X_{it}$ given $z_{st}$:
$Q_{t}(a_{jk}\mid z_{st})\geq\Pr\left(Y_{it}\in K,a_{jk}\geq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z_{st}\right).$
Similarly, an upper bound for the conditional probability $Q_{t}(a_{jk}\mid w_{st})$ is derived as follows:
$Q_{t}(a_{jk}\mid w_{st})\leq\Pr(Y_{it}\in K\mid w_{st})\mathbf{\mathbbm{1}}\{a_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,k\in K^{c}\}+\\ 1-\mathbf{\mathbbm{1}}\{a_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,k\in K^{c}\}.$
The above inequality holds since either $a_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ and the probability $Q_{t}(a_{jk}\mid w_{st})$ is weakly increasing or the upper bound is one.
After taking expectation of $X_{ist}$ given $z_{st}$, the upper bound for $Q_{t}(a_{jk}\mid z_{st})$ is obtained as
$Q_{t}(a_{jk}\mid z_{st})\leq\Pr\left(Y_{it}\in K,a_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z_{st}\right)\\ +1-\Pr\left(a_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z_{st}\right).$
Rearranging the above formula yields
$Q_{t}(a_{jk}\mid z_{st})\leq 1-\Pr\left(Y_{it}\in K^{c},a_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z_{st}\right).$
Under Assumption 1, the conditional probability $Q_{t}(a_{jk}\mid z_{st})$ is the same over the two periods $(s,t)$. Therefore, the upper bound at period $s$ of $Q_{t}(a_{jk}\mid z_{st})$ should be larger than the lower bound at period $t$, yielding the identifying condition (10) as follows:
$1-\Pr(Y_{is}\in K^{c},(z_{js}-z_{ks})^{\prime}\beta_{0}+(X_{ijs}-X_{iks})^{\prime}\gamma_{0}\geq a_{jk}\ \forall j\in K,k\in K^{c}\mid z_{st})\\ \geq\Pr(Y_{it}\in K,(z_{jt}-z_{kt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\leq a_{jk}\ \forall j\in K,k\in K^{c}\mid z_{st}).$
A.9 Proof of Proposition 8
Proof.
@@ A.9 Proof of Proposition 8
Since the observed outcome $Y_{it}$ is censored, we either observe $Y_{it}=y>0$ or $Y_{it}=0$. Let $v_{it}=-(\alpha_{i}+\epsilon_{it})$, the conditional probability of $Y_{it}=0$ is given as,
$\Pr(Y_{it}=0\mid w_{st})=\Pr(Y^{*}_{it}\leq 0\mid w_{st})=\Pr(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\mid w_{st}).$
@@
$\Pr(Y_{it}\leq y\mid w_{st})$ $=\Pr(Y^{*}_{it}\leq 0,Y_{it}\leq y\mid w_{st})+\Pr(0<Y^{*}_{it},Y_{it}\leq y\mid w_{st})$
$=\Pr(Y^{*}_{it}\leq 0\mid w_{st})+\Pr(0<Y^{*}_{it}\leq y\mid w_{st})$
$=\Pr(Y^{*}_{it}\leq y\mid w_{st})$
$=\Pr(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-y\mid w_{st}).$
Combining the two scenarios, the full conditional distributional of $Y_{it}\mid W_{ist}$ is characterized as follows:
@@
$\Pr(Y_{it}\leq y\mid w_{st})=\left\{\begin{aligned} \Pr(v_{it}\geq z_{t}^{\prime}\beta_{0}+&x_{t}^{\prime}\gamma_{0}-y\mid w_{st})\quad&\text{ if }y\geq 0,\\ &0\quad&\text{ if }y<0.\end{aligned}\right.$
Given observed distribution of $Y_{it}\mid W_{ist}$, the identification approach is to bound the distribution $\Pr(v_{it}\leq a\mid z_{st})$. We first look at the upper bound for the conditional distribution $v_{it}\mid w_{st}$ given all covariates $W_{ist}=w_{st}$.
$\Pr(v_{it}\geq a\mid w_{st})\leq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a)\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\geq a\}+\\ \Pr(Y_{it}=0\mid z_{st})\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<a\},$
@@
where the above condition holds since either $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a\geq 0$ so that there exists $y=z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a\geq 0$ such that $\Pr(Y_{it}\leq y\mid w_{st})=\Pr(v_{it}\geq a\mid w_{st})$, or $\Pr(v_{it}\geq a\mid w_{st})\leq\Pr(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\mid w_{st})=\Pr(Y_{it}=0\mid w_{st})$ when $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<a$.
@@
Taking expectation over the endogenous covariate $X_{ist}$ yields the upper bound for the distribution $v_{it}\mid z_{st}$:
$\Pr(v_{it}\geq a\mid z_{st})\leq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-a,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st})+\\ \Pr(Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}<a\mid z_{st}).$
The above upper bound can be also also expressed as
$\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-a,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st})+\Pr(Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}<a\mid z_{st})$
$=$ $\Pr(0<Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-a,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\geq a\mid z_{st})+\Pr(Y_{it}=0\mid z_{st})$
$=$ $\Pr(0<Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-a\mid z_{st})+\Pr(Y_{it}=0\mid z_{st}).$
Similarly, the conditional distribution $v_{it}\mid w_{st}$ can be bounded below
$\Pr(v_{it}\geq a\mid w_{st})\geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a),$
@@
where the above condition holds since either $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a\geq 0$ so that there exists $y=z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a\geq 0$ such that $\Pr(Y_{it}\leq y\mid w_{st})=\Pr(v_{it}\geq a\mid w_{st})$, or the lower bound is zero when $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<a$.
Taking expectation over $X_{ist}$ leads to the following lower bound:
$\Pr(v_{it}\geq a\mid z_{st})\geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-a\mid z_{st}).$
The conditional stationarity condition requires $v_{is}\mid Z_{ist}\stackrel{{\scriptstyle d}}{{\sim}}v_{it}\mid Z_{ist}$, which implies that the bounds for $v_{it}\mid Z_{ist}$ must have intersections over any pair of periods $(s,t)$. This restriction generates the following identifying condition for $\theta_{0}$:
$\Pr(0<Y_{is}\leq z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma-a\mid z_{st})+\Pr(Y_{is}=0\mid z_{st})\geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-a\mid z_{st})$
for any $a\in\mathcal{R}$, any $s,t$, and any $z_{st}$.
A.10 Dynamic Censored Models with Latent Dependent Variables

### v3

@@ A.7 Proof of Proposition 9
Since the observed outcome $Y_{it}$ is censored at 0, we either observe $Y_{it}=y>0$ or $Y_{it}=0$. Let $v_{it}:=-(\alpha_{i}+\epsilon_{it})$, the conditional probability of $Y_{it}=0$ is given as,
$\mathbb{P}(Y_{it}=0\mid w)=\mathbb{P}(Y_{it}^{*}\leq 0\mid w)=\mathbb{P}(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\mid w).$
@@
$\mathbb{P}(Y_{it}\leq y\mid w)$ $=\mathbb{P}(Y_{it}^{*}\leq 0,\ Y_{it}\leq y\mid w)+\mathbb{P}(0<Y_{it}^{*},\ Y_{it}\leq y\mid w)$
$=\mathbb{P}(Y_{it}^{*}\leq 0\mid w)+\mathbb{P}(0<Y_{it}^{*}\leq y\mid w)$
$=\mathbb{P}(Y_{it}^{*}\leq y\mid w)$
$=\mathbb{P}(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-y\mid w).$
Combining the two scenarios, the conditional distributional of $Y_{it}\mid W_{i}$ is characterized as follows:
@@
$\mathbb{P}(Y_{it}\leq y\mid w)=\left\{\begin{aligned} \mathbb{P}(v_{it}\geq z_{t}^{\prime}\beta_{0}+&x_{t}^{\prime}\gamma_{0}-y\mid w)\quad&\text{ if }y\geq 0,\\ &0\quad&\text{ if }y<0.\end{aligned}\right.$
Given observed distribution of $Y_{it}\mid W_{i}$, we can bound the distribution $\mathbb{P}(v_{it}\leq c\mid w)$ above as
$\mathbb{P}(v_{it}\geq c\mid w)\leq\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-c\mid w)\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\geq c\}+\\ \mathbb{P}(Y_{it}=0\mid w)\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<c\},$
@@
where the above condition holds since either $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-c\geq 0$ so that there exists $y=z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-c\geq 0$ such that $\mathbb{P}(Y_{it}\leq y\mid w)=\mathbb{P}(v_{it}\geq c\mid w)$, or $\mathbb{P}(v_{it}\geq c\mid w)\leq\mathbb{P}(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\mid w)=\mathbb{P}(Y_{it}=0\mid w)$ when $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<c$.
@@
Taking expectation over the endogenous covariate $X_{i}$ yields the upper bound for the distribution $v_{it}\mid Z_{i}=z$:
$\mathbb{P}(v_{it}\geq c\mid z)\leq\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-c,\ z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq c\mid z)+\\ \mathbb{P}(Y_{it}=0,\ z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}<c\mid z).$
Rearranging the formula, the above upper bound can be equivalently written as
$\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-c,\ z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq c\mid z)+\mathbb{P}(Y_{it}=0,\ z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}<c\mid z)$
$=$ $\mathbb{P}(0<Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-c,\ z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\geq c\mid z)+\mathbb{P}(Y_{it}=0\mid z)$
$=$ $\mathbb{P}(0<Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-c\mid z)+\mathbb{P}(Y_{it}=0\mid z).$
Similarly, the conditional distribution $v_{it}\mid w$ can be bounded below
$\mathbb{P}(v_{it}\geq c\mid w)\geq\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-c\mid w),$
@@
where the above condition holds since either $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-c\geq 0$ so that there exists $y=z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-c\geq 0$ such that $\mathbb{P}(Y_{it}\leq y\mid w)=\mathbb{P}(v_{it}\geq c\mid w)$, or the lower bound is zero when $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<c$.
Taking expectation over $X_{i}$ leads to the following lower bound:
$\mathbb{P}(v_{it}\geq c\mid z)\geq\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-c\mid z).$
Given the bounds on the distribution $\mathbb{P}(v_{it}\geq c\mid z)$, the partial stationarity assumption implies the following identifying restriction for $\theta_{0}$:
$\max_{t}\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-c\mid z)\leq\max_{s}\{\mathbb{P}(0<Y_{is}\leq z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma-c\mid z)+\mathbb{P}(Y_{is}=0\mid z)\},$
for any $c\in\mathcal{R}$ and any $z$. ∎
A.8 Dynamic Censored Models with Latent Dependent Variables